# On the Equivalence of Random Network Distillation, Deep Ensembles, and Bayesian Inference

### v1

@@ 4 Equivalence of Random Network Distillation & Bayesian Posteriors
Proof sketch. The result follows by combining Theorem 3 and Proposition 4, provided that the GP governing the predictor initialization $\kappa^{u}_{xx^{\prime}}$ and the target function $\kappa^{\tilde{g}}_{xx^{\prime}}$ are independent. Owing to the fact that the parameters $\vartheta_{0}$ and $\psi_{0}$ are drawn independently, the independence between $u(x;\vartheta_{0})$ and $\tilde{g}(x;\vartheta_{0},\psi_{0})$ is apparent by rewriting the covariance $\mathbb{E}[u(x;\vartheta_{0})\tilde{g}(x;\vartheta_{0},\psi_{0})]$ in terms of conditional expectations on $\vartheta_{0}$ by the law of total expectation. Furthermore, since $\Theta_{xx^{\prime}}=\Theta^{L}_{xx^{\prime}}+\Theta^{\leq L-1}_{xx^{\prime}}$ and $\kappa^{\tilde{g}}_{xx^{\prime}}=\Theta^{\leq L-1}_{xx^{\prime}}$, $\kappa^{u}_{xx^{\prime}}=\Theta^{L}_{xx^{\prime}}$, we have that $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\Theta_{xx^{\prime}}$. In other words, the GP kernel of initial errors aligns with the NTK of the online predictor, such that the distribution of post-convergence errors in Theorem 3 simplifies significantly. This same covariance function indeed also defines the posterior predictive distribution of infinitely wide neural networks as described by the GP with prior $\mathcal{GP}(0,\Theta_{xx^{\prime}})$ and conditioned on $(\mathcal{X},\mathcal{Y})$.
@@
Theorem 4 shows that with a specifically engineered target function, the RND error signal $\epsilon^{b}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})=u(x;\vartheta_{\infty})-\tilde{g}(x;\vartheta_{0},\psi_{0})$ is no longer just related to ensemble variance, but rather becomes a direct sample from the centered posterior predictive distribution of a Bayesian model whose prior kernel is the NTK itself. This novel result provides a direct bridge between RND and Bayesian inference in the limit of infinite network width, providing a useful insight: the error signal generated by this modified RND procedure is not merely a heuristic measure of distance, but is itself a random draw from the (centered) Bayesian posterior predictive distribution of an NTK-based GP. This direct distributional equivalence has immediate practical implications, for example prescribing rather straightforwardly how this Bayesian form of RND can be used for exact posterior sampling. By applying Proposition 3.1 to the multi-headed Bayesian RND architecture [^6] In a multi-headed architecture, the Bayesian target function described in Proposition 4 becomes a JVP. Several common machine learning libraries (e.g., JAX [jax2018github] offer dedicated algorithms to compute such JVPs efficiently., in contrast to obtaining samples from deep ensembles as done in Theorem 3.1, we now obtain several independent samples from the centered posterior predictive distribution through $\epsilon^{b}_{i}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})=u_{i}(x;\vartheta_{\infty})-\tilde{g}_{i}(x;\vartheta_{0},\psi_{0})$. The below corollary details how this can be leveraged to conduct a posterior sampling procedure, requiring access only to a mean estimate and a single Bayesian RND model.
@@
Let $\mathcal{N}\bigl(\mu^{b}(x)\,,\,\,\Sigma^{b}_{xx^{\prime}}\bigr)$ be the posterior predictive distribution of an infinitely wide neural network conditioned on $x$ with mean $\mu^{b}(x)=\Theta_{x\mathcal{X}}\Theta^{-1}_{\mathcal{X}\mathcal{X}}\mathcal{Y}$ and covariance $\Sigma^{b}_{xx^{\prime}}=\Theta_{xx^{\prime}}-\Theta_{x\mathcal{X}}\Theta^{-1}_{\mathcal{X}\mathcal{X}}\Theta_{\mathcal{X}x^{\prime}}$. Suppose $\tilde{\mu}(x;\theta_{\infty})\approx\mu^{b}(x)$ is an estimate of the mean function and let $\{\epsilon^{b}_{i}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})\}_{i=1}^{K}$ be error functions of a $K$-head Bayesian RND model as defined in Theorem 4.
@@
Proof sketch. The result follows directly from Theorem (4) and application of the independence argument of Proposition (3.1) to the multi-headed setting.
@@ 5 Numerical Analysis
We proceed with a numerical analysis to validate the thus far presented results. In the following, we study how predictive RND errors of the Bayesian RND model correlate with draws from a Bayesian deep ensemble [heBayesianDeepEnsembles2020a] in practice. To this end, we train both models with a single-layer fully connected neural network with SiLU activations [elfwing2018sigmoid] on a synthetic dataset with training input elements $\{x_{train}^{1},...,x_{train}^{N}\}\sim\mathcal{N}(0,I)$, $x_{i}\in\mathbb{R}^{3}$ and equally for a test set $\{x_{test}^{1},...,x_{test}^{\tilde{N}}\}$. Ensemble models are fit to a toy target function, and Bayesian RND models optimized as described above. The variance of the true underlying GP is approximated with Monte-Carlo estimates of 1024 independent models and a single Bayesian RND model with 1024 heads, such that a small residual amount of discrepancy is to be expected. Fig. 1 shows test evaluations of the predictive variances of a Bayesian deep ensemble compared to squared RND errors (per test sample), which become more strongly correlated and correctly calibrated in scale at higher network width.
Figure 1: Comparison of test-set evaluations of predictive variances by Bayesian ensembles and Bayesian RND.
@@ B.1 Ensemble Equivalence
{restatable}
thmrndwidennconverged[jacotNeuralTangentKernel2020](Post-convergence neural network function) In the limit of infinite layer widths $n\xrightarrow{}\infty$ and infinite time $t\xrightarrow{}\infty$, the output function of a neural network $f(x;\theta_{\infty})$ with NTK parametrization according to Eq. 1 is given by
@@
where we used the shorthand $\Theta_{xx^{\prime}}\equiv\Theta(x,x^{\prime})$. Proof sketch. By taking the infinite width limit $n\to\infty$, we obtain a linear ODE from Eq. (2). Through an exponential ansatz, its explicit solution with initial condition $f(x;\theta_{0})$ is given by $f(x;\theta_{t})=f(x;\theta_{0})+\Theta_{x\mathcal{X}}\Theta_{\mathcal{X}\mathcal{X}}^{-1}(I-e^{-t\Theta_{\mathcal{X}\mathcal{X}}})(\mathcal{Y}-f(\mathcal{X};\theta_{0}))$ . Assuming the training Gram matrix $\Theta_{\mathcal{X}\mathcal{X}}$ is positive definite (and thus invertible), the exponential term decays to zero as $t\to\infty$, yielding the kernel regression formula in Proposition (B.1). See jacotNeuralTangentKernel2020 and Appendix B.1.1.
@@
B.1.2 Proof of Theorem 2
[…]
Proof sketch. We use the fact that $f(x;\theta_{\infty})$ can be written as a linear combination of the test initialization $f(x;\theta_{0})$ and the training initialization $f(\mathcal{X};\theta_{0})$. Both these identities are described probabilistically by the NNGP $f(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{xx^{\prime}})$, and $f(\mathcal{X};\theta_{0})\sim\mathcal{GP}(0,\kappa_{\mathcal{X}\mathcal{X}})$. Applying a linear transformation to a GP yields another GP [rasmussen2006gp], meaning $f(x;\theta_{\infty})$ also follows a GP. Propagating the prior covariance $\kappa$ through the linear transformation described by Proposition B.1 reveals the expression for the post-convergence covariance function $\Sigma^{f}_{\mathcal{X}_{T}\mathcal{X}_{T}}(\theta_{\infty})$ given in Theorem 2.
@@ B.1.2 Proof of Theorem 2
B.1.3 Proof of Theorem 3
[…]
@@ B.1.3 Proof of Theorem 3
{restatable}
prprndindependencenngp[lee2017deep] (Recursive NNGP formulation) At initialization $t=0$ and in the limit $n\to\infty$, the $i$-th output at layer $l$, $z_{i}^{l}(x;\theta_{0}^{\leq l})$, converges to a GP with zero mean and covariance function $\kappa_{ii}^{l}(x,x^{\prime})$ given by
@@
{restatable}
prprndindependencentk[jacotNeuralTangentKernel2020] (Recursive NTK formulation) In the limit $n\to\infty$, the neural tangent kernel $\Theta^{l}_{ii}(x,x^{\prime})$ of the $i$-th output $z_{i}^{l}(x;\theta_{0}^{\leq l})$ at layer $l$, defined as the gradient inner product
@@
Note that the above $r.h.s$ involves gradients w.r.t. last-layer parameters, i.e. the post-activation outputs of the previous layer, and by the same arguments as in the NNGP derivation of Proposition B.1.4, this is a sum of independent post activations s.t. in the limit $n_{l-1}\xrightarrow{}\infty$
@@
We begin by deriving the training dynamics for the output $f_{i}(x;\theta_{t})$ analogously to the proof of Proposition B.1. We denote by $\mathcal{Y}_{i}$ the labels used to train the function $f_{i}(x;\theta_{t})$. By Proposition B.1.4, the training dynamics of $f_{i}(x;\theta_{t})$ and $f_{j}(x;\theta_{t})$ are decoupled for $i\neq j$ and we can thus derive Eq. 23 analogously for individual output heads $i$. Taking the infinite width limit, we obtain at time $t$
@@
Thus, the output head $f_{i}(x;\theta_{t})$ at time $t$ is a deterministic function of its own initialization only, which itself is characterized by a GP $f_{i}(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{ii}(x,x^{\prime}))$ that is independent of output heads $j\neq i$ by Proposition B.1.4. And thus, since $f_{i}(x;\theta_{t})$ is an affine transform of its own independent initialization terms $f_{i}(x;\theta_{0})$ and $f_{i}(\mathcal{X};\theta_{0})$, it too must follow an independent GP with $\mathbb{E}_{\theta_{0}}[f_{i}(x;\theta_{t})f_{i}(x^{\prime};\theta_{t})]=\Sigma(x,x^{\prime};\theta_{t})$ and in particular $\mathbb{E}_{\theta_{0}}[f_{i}(x;\theta_{t})f_{j}(x^{\prime};\theta_{t})]=0$ if $i\neq j$. ∎
@@
The proof follows by combining the results of Propositions (3) and (3.1). We define a multiheaded RND predictor with $K$ output heads $\{u_{i}(x,\vartheta_{t})\}_{i=1}^{K}$ and a fixed multiheaded target network $\{g_{i}(x_{t};\psi_{0})\}_{i=1}^{K}$ of equivalent architecture as $u_{i}$ (i.e., both corresponding to the same NTK $\Theta$) with the corresponding prediction errors $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ accordingly. Let $u_{i}(x,\vartheta_{t})$ be trained such that each head $i$ is trained to match the $i$-th target output $g_{i}(x;\psi_{0})$.
@@
As previously established in the proof of Proposition 3.1, the multi-headed functions $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ follow equivalent learning dynamics as their scalar-output counterparts. The post-convergence distribution of individual heads $\epsilon_{i}(x;\vartheta_{\infty},\psi_{0})$ must therefore equal the scalar-output post-convergence distribution established in Theorem 3. Consequently, the errors $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ are independent and identically distributed draws from a Gaussian with mean and covariance
@@
where $\Sigma^{\epsilon}_{xx}(\vartheta_{\infty},\psi_{0})$ is the variance of the GP described in Theorem 3.
Conversely, a set of $K+1$ independent neural networks arranged to a deep ensemble $\{f(x;\theta_{\infty}^{i})\}_{i=1}^{K+1}$ in the infinite width limit $n\to\infty$ and at convergence $t\to\infty$ are by definition i.i.d. samples from the GP described in Theorem 2. As before, the empirical variance defined as $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})=\frac{1}{K}\sum_{i=1}^{K+1}\bigl(f(x;\theta^{i}_{\infty})-\frac{1}{K+1}\sum_{j=1}^{K+1}f(x;\theta^{j}_{\infty})\bigr)^{2}$ distributes as a scaled Chi-squared distribution with $K$ degrees of freedom
@@
where $\Sigma^{f}_{xx}(\theta_{\infty})$ is the variance of the GP described in Theorem 2.

### v2

@@ 4 Equivalence of Random Network Distillation & Bayesian Posteriors
Proof sketch. The result follows by combining Theorem 3.1 and Proposition 4.1, provided that the GP governing the predictor initialization $\kappa^{u}_{xx^{\prime}}$ and the target function $\kappa^{\tilde{g}}_{xx^{\prime}}$ are independent. Owing to the fact that the parameters $\vartheta_{0}$ and $\psi_{0}$ are drawn independently, the independence between $u(x;\vartheta_{0})$ and $\tilde{g}(x;\vartheta_{0},\psi_{0})$ is apparent by rewriting the covariance $\mathbb{E}[u(x;\vartheta_{0})\tilde{g}(x;\vartheta_{0},\psi_{0})]$ in terms of conditional expectations on $\vartheta_{0}$ by the law of total expectation. Furthermore, since $\Theta_{xx^{\prime}}=\Theta^{L}_{xx^{\prime}}+\Theta^{\leq L-1}_{xx^{\prime}}$ and $\kappa^{\tilde{g}}_{xx^{\prime}}=\Theta^{\leq L-1}_{xx^{\prime}}$, $\kappa^{u}_{xx^{\prime}}=\Theta^{L}_{xx^{\prime}}$, we have that $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\Theta_{xx^{\prime}}$. In other words, the GP kernel of initial errors aligns with the NTK of the online predictor, such that the distribution of post-convergence errors in Theorem 3.1 simplifies significantly. This same covariance function indeed also defines the posterior predictive distribution of infinitely wide neural networks as described by the GP with prior $\mathcal{GP}(0,\Theta_{xx^{\prime}})$ and conditioned on $(\mathcal{X},\mathcal{Y})$.
@@
Theorem 4.2 shows that with a specifically engineered target function, the RND error signal $\epsilon^{b}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})=u(x;\vartheta_{\infty})-\tilde{g}(x;\vartheta_{0},\psi_{0})$ is no longer just related to ensemble variance, but rather becomes a direct sample from the centered posterior predictive distribution of a Bayesian model whose prior kernel is the NTK itself. This novel result provides a direct bridge between RND and Bayesian inference in the limit of infinite network width, providing a useful insight: the error signal generated by this modified RND procedure is not merely a heuristic measure of distance, but is itself a random draw from the (centered) Bayesian posterior predictive distribution of an NTK-based GP. This direct distributional equivalence has immediate practical implications, for example prescribing rather straightforwardly how this Bayesian form of RND can be used for exact posterior sampling. By applying Proposition 3.3 to the multi-headed Bayesian RND architecture [^6] In a multi-headed architecture, the Bayesian target function described in Proposition 4.2 becomes a JVP. Several common machine learning libraries (e.g., JAX [jax2018github] offer dedicated algorithms to compute such JVPs efficiently., in contrast to obtaining samples from deep ensembles as done in Theorem 3.4, we now obtain several independent samples from the centered posterior predictive distribution through $\epsilon^{b}_{i}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})=u_{i}(x;\vartheta_{\infty})-\tilde{g}_{i}(x;\vartheta_{0},\psi_{0})$. The below corollary details how this can be leveraged to conduct a posterior sampling procedure, requiring access only to a mean estimate and a single Bayesian RND model.
@@
Let $\mathcal{N}\bigl(\mu^{b}(x)\,,\,\,\Sigma^{b}_{xx^{\prime}}\bigr)$ be the posterior predictive distribution of an infinitely wide neural network conditioned on $x$ with mean $\mu^{b}(x)=\Theta_{x\mathcal{X}}\Theta^{-1}_{\mathcal{X}\mathcal{X}}\mathcal{Y}$ and covariance $\Sigma^{b}_{xx^{\prime}}=\Theta_{xx^{\prime}}-\Theta_{x\mathcal{X}}\Theta^{-1}_{\mathcal{X}\mathcal{X}}\Theta_{\mathcal{X}x^{\prime}}$. Suppose $\tilde{\mu}(x;\theta_{\infty})\approx\mu^{b}(x)$ is an estimate of the mean function and let $\{\epsilon^{b}_{i}(x;\vartheta_{\infty},\vartheta_{0},\psi_{0})\}_{i=1}^{K}$ be error functions of a $K$-head Bayesian RND model as defined in Theorem 4.2.
@@
Proof sketch. The result follows directly from Theorem (4.2) and application of the independence argument of Proposition (3.3) to the multi-headed setting.
@@ 5 Numerical Analysis
We proceed with a numerical analysis to validate the thus far presented results. In the following, we study how predictive RND errors relate to predictive variances of deep ensembles in practice, both in the standard and Bayesian settings. To this end, we train two-layer connected neural networks with SiLU activations [elfwing2018sigmoid] on a synthetic dataset with $N=10$ train and $\tilde{N}=5000$ test samples from an isotropic Gaussian $x_{i}\sim\mathcal{N}(0,I_{3})$. Ensemble models are fit to a toy target function, and multiheaded RND models optimized as described above. The variance of the true underlying GP is approximated with Monte-Carlo estimates of 512 independent models and a single Bayesian RND model with 512 heads, such that a small residual amount of discrepancy is to be expected. Fig. 1 shows a stark decrease in average squared discrepancy between test evaluations of predictive ensemble variances and RND errors as model width increases, a trend in line with our theoretical derivations and present even at practical network widths. Further evaluations and details of this experiment are reported in Appendix C.
Figure 1: Test-set errors between predictive variances of (Bayesian) ensembles and self-predictive errors of (Bayesian) RND vanish with large layer widths.
@@ B.1 Ensemble Equivalence
Theorem B.1.
[jacotNeuralTangentKernel2020](Post-convergence neural network function) In the limit of infinite layer widths $n\xrightarrow{}\infty$ and infinite time $t\xrightarrow{}\infty$, the output function of a neural network $f(x;\theta_{\infty})$ with NTK parametrization according to Eq. 1 is given by
@@
where we used the shorthand $\Theta_{xx^{\prime}}\equiv\Theta(x,x^{\prime})$.
Proof sketch. By taking the infinite width limit $n\to\infty$, we obtain a linear ODE from Eq. (2). Through an exponential ansatz, its explicit solution with initial condition $f(x;\theta_{0})$ is given by $f(x;\theta_{t})=f(x;\theta_{0})+\Theta_{x\mathcal{X}}\Theta_{\mathcal{X}\mathcal{X}}^{-1}(I-e^{-t\Theta_{\mathcal{X}\mathcal{X}}})(\mathcal{Y}-f(\mathcal{X};\theta_{0})).$ Assuming the training Gram matrix $\Theta_{\mathcal{X}\mathcal{X}}$ is positive definite (and thus invertible), the exponential term decays to zero as $t\to\infty$, yielding the kernel regression formula in Proposition (B.1). See jacotNeuralTangentKernel2020 and Appendix B.1.1.
@@
B.1.2 Proof of Theorem 2.1
We restate Theorem 2.1 for convenience. See 2.1
Proof sketch. We use the fact that $f(x;\theta_{\infty})$ can be written as a linear combination of the test initialization $f(x;\theta_{0})$ and the training initialization $f(\mathcal{X};\theta_{0})$. Both these identities are described probabilistically by the NNGP $f(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{xx^{\prime}})$, and $f(\mathcal{X};\theta_{0})\sim\mathcal{GP}(0,\kappa_{\mathcal{X}\mathcal{X}})$. Applying a linear transformation to a GP yields another GP [rasmussen2006gp], meaning $f(x;\theta_{\infty})$ also follows a GP. Propagating the prior covariance $\kappa$ through the linear transformation described by Proposition B.1 reveals the expression for the post-convergence covariance function $\Sigma^{f}_{\mathcal{X}_{T}\mathcal{X}_{T}}(\theta_{\infty})$ given in Theorem 2.1.
@@
B.1.3 Proof of Theorem 3.1
We restate Theorem 3.1 for convenience. See 3.1
@@
Proposition B.2.
[lee2017deep] (Recursive NNGP formulation) At initialization $t=0$ and in the limit $n\to\infty$, the $i$-th output at layer $l$, $z_{i}^{l}(x;\theta_{0}^{\leq l})$, converges to a GP with zero mean and covariance function $\kappa_{ii}^{l}(x,x^{\prime})$ given by
@@
Proposition B.3.
[jacotNeuralTangentKernel2020] (Recursive NTK formulation) In the limit $n\to\infty$, the neural tangent kernel $\Theta^{l}_{ii}(x,x^{\prime})$ of the $i$-th output $z_{i}^{l}(x;\theta_{0}^{\leq l})$ at layer $l$, defined as the gradient inner product
@@
Note that the above $r.h.s$ involves gradients w.r.t. last-layer parameters, i.e. the post-activation outputs of the previous layer, and by the same arguments as in the NNGP derivation of Proposition B.2, this is a sum of independent post activations s.t. in the limit $n_{l-1}\xrightarrow{}\infty$
@@
We begin by deriving the training dynamics for the output $f_{i}(x;\theta_{t})$ analogously to the proof of Proposition B.1. We denote by $\mathcal{Y}_{i}$ the labels used to train the function $f_{i}(x;\theta_{t})$. By Proposition B.3, the training dynamics of $f_{i}(x;\theta_{t})$ and $f_{j}(x;\theta_{t})$ are decoupled for $i\neq j$ and we can thus derive Eq. 23 analogously for individual output heads $i$. Taking the infinite width limit, we obtain at time $t$
@@
Thus, the output head $f_{i}(x;\theta_{t})$ at time $t$ is a deterministic function of its own initialization only, which itself is characterized by a GP $f_{i}(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{ii}(x,x^{\prime}))$ that is independent of output heads $j\neq i$ by Proposition B.2. And thus, since $f_{i}(x;\theta_{t})$ is an affine transform of its own independent initialization terms $f_{i}(x;\theta_{0})$ and $f_{i}(\mathcal{X};\theta_{0})$, it too must follow an independent GP with $\mathbb{E}_{\theta_{0}}[f_{i}(x;\theta_{t})f_{i}(x^{\prime};\theta_{t})]=\Sigma(x,x^{\prime};\theta_{t})$ and in particular $\mathbb{E}_{\theta_{0}}[f_{i}(x;\theta_{t})f_{j}(x^{\prime};\theta_{t})]=0$ if $i\neq j$. ∎
@@
The proof follows by combining the results of Propositions (3.1) and (3.3). We define a multiheaded RND predictor with $K$ output heads $\{u_{i}(x,\vartheta_{t})\}_{i=1}^{K}$ and a fixed multiheaded target network $\{g_{i}(x_{t};\psi_{0})\}_{i=1}^{K}$ of equivalent architecture as $u_{i}$ (i.e., both corresponding to the same NTK $\Theta$) with the corresponding prediction errors $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ accordingly. Let $u_{i}(x,\vartheta_{t})$ be trained such that each head $i$ is trained to match the $i$-th target output $g_{i}(x;\psi_{0})$.
@@
As previously established in the proof of Proposition 3.3, the multi-headed functions $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ follow equivalent learning dynamics as their scalar-output counterparts. The post-convergence distribution of individual heads $\epsilon_{i}(x;\vartheta_{\infty},\psi_{0})$ must therefore equal the scalar-output post-convergence distribution established in Theorem 3.1. Consequently, the errors $\{\epsilon_{i}(x;\vartheta_{t},\psi_{0})\}_{i=1}^{K}$ are independent and identically distributed draws from a Gaussian with mean and covariance
@@
where $\Sigma^{\epsilon}_{xx}(\vartheta_{\infty},\psi_{0})$ is the variance of the GP described in Theorem 3.1.
Conversely, a set of $K+1$ independent neural networks arranged to a deep ensemble $\{f(x;\theta_{\infty}^{i})\}_{i=1}^{K+1}$ in the infinite width limit $n\to\infty$ and at convergence $t\to\infty$ are by definition i.i.d. samples from the GP described in Theorem 2.1. As before, the empirical variance defined as $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})=\frac{1}{K}\sum_{i=1}^{K+1}\bigl(f(x;\theta^{i}_{\infty})-\frac{1}{K+1}\sum_{j=1}^{K+1}f(x;\theta^{j}_{\infty})\bigr)^{2}$ distributes as a scaled Chi-squared distribution with $K$ degrees of freedom
@@
where $\Sigma^{f}_{xx}(\theta_{\infty})$ is the variance of the GP described in Theorem 2.1.